OpenStudy (anonymous):

Find the distance D between the point (1, 4, 2) and the line passing through the points (−1, 3, −1) and (1, 2, 3).

OpenStudy (anonymous):

it says the answer is sqrt(23/7) but im not sure how it got that.

OpenStudy (phi):

Here is one way: call pt (-1,3,-1) p0 and pt (1,2,3) p1 and pt (1,4,2) p2 subtract p0 from the other two points: u= p1-p0 and q= p2-p0 this will give you two vectors with a common origin normalize the vector u to unit length: u1=u/ | u | do a dot product between u1 and q: u1 dot q = x now find y: y= q- x*u1 y is perpendicular to u, and its length is the distance between point p2 and the line |y| is the answer.

OpenStudy (phi):

The good news is that procedure got 1.8127 which matches sqrt(23/7) The bad news is I did not do it by hand, I don't know the details...

OpenStudy (phi):

Here are some of the numbers p2 = 1 4 2 p0 = -1 3 -1 p1 = 1 2 3 q = 2 1 3 u = 2 -1 4 length of u sqrt(21) u1= u/sqrt(21) b= dot(p,u1) b =3.2733 x= b*u1 x = 1.4286 -0.7143 2.8571 y= q- x y = 0.5714 1.7143 0.1429 length(y)= 1.8127